OpenStudy (amtran_bus):

Solve the system for c1 and c2

OpenStudy (amtran_bus):

c1+c2=0 c1+2c2=-1

OpenStudy (anonymous):

multiply equation 1 by -1. then add them

OpenStudy (amtran_bus):

Ok. So that is -c1-c2=0 c1+2c2=-1 Or c2= -1

OpenStudy (amtran_bus):

-c1+1=0 so c1 = -1?

OpenStudy (amtran_bus):

Nope, c1 = 1

OpenStudy (emmatassone):

good

OpenStudy (amtran_bus):

Thanks all